OpenStudy (anonymous):

I'm confused. x + 4 – 7x = 22 A. 3 B. –3 C. 5 D. –5

OpenStudy (bibby):

combine the like terms

OpenStudy (anonymous):

@bibby it was asking me to plug in the numbers to see which one equals 22. Bcz it told me to solve.

OpenStudy (bibby):

before you plug anything in, combine the like terms. that means the x's and numbers. then solve for x x + 4 – 7x = 22 x-7x +4=22 combine the x's -6x+4=22 subtract 4 from both sides -6x=18

OpenStudy (anonymous):

ok thanks @bibby now i understand.

OpenStudy (anonymous):

Are you in 7th grade?

OpenStudy (anonymous):

@Tomboyz no 8th